# Parametricity via Cohesion

### v2

@@ 3.2 Cohesion & Coherence
As we have seen, there appears to be an intimate link between parametricity, cohesion, and coherence, demonstrated (e.g.) by the above-given proof of induction for $S^{1}$. The existence of such a link is further supported by other recent developments in HoTT and related fields. E.g. Cavallo & Harper [CH21] have used their system of internal parametricity for Cubical Type Theory to derive coherence theorems for the smash product. More recently still, Kolomatskaia & Shulman [KS24] have introduced their system of Displayed Type Theory, which utilizes yet another form of internal parametricity to solve the problem of representing semi-simplicial types in HoTT.
In outlining the above framework for parametricity in terms of cohesion, I hope to have taken a first step toward the unification of these various systems that use parametricity to tackle instances of the coherence problem. Cavallo & Harper’s system is readily assimilated to this framework, as are other related systems that take a cubical approach to internal parametricity, such as that of Nuyts, Devriese & Vezzossi [NVD17], since these all take their semantics in various topoi of bicubical sets, which all are sufficiently cohesive over corresponding topoi of cubical sets. On the other hand, Kolomatskaia & Shulman’s system cannot be assimilated in quite the same way, since their system takes its semantics in the $\infty$-topos of augmented semisimplicial spaces, which is not cohesive over spaces. It thus remains to be seen if the above framework can be generalized so as to be inclusive of this example as well. Alternatively, one might seek to generalize or modify Kolomatskaia & Shulman’s system so as to be interpretable in an $\infty$-topos that is cohesive over spaces, e.g. the $\infty$-topos of simplicial spaces. If this latter proves feasible, then this in turn would reveal yet further connections between parametricity via cohesion and another prominent attempt at a solution to the coherence problem, namely Riehl & Shulman’s Simplicial Type Theory [RS23], which takes its semantics in simplicial spaces.
@@
3.3 Acknowledgements
The origins of this paper trace to research I did as an undergraduate at Merton College, Oxford, in the Summer of 2020, as part of the Merton College Summer Projects Scheme. Naturally, this work was conducted at a time of considerable stress for myself and the world at large, and I am massively indebted to the academic support staff at Merton, who were immensely helpful in keeping me afloat at that time. I am particularly grateful to Katy Fifield, Jemma Underdown, and Jane Gover for the support they provided to me in the course of my undergraduate studies. More recently, I am thankful to Frank Pfenning and Steve Awodey for their encouragement in continuing to pursue this line of research, and to Tesla Zhang for many helpful tips on how to make the best use of Agda for this paper.

### v3

@@ 3.2 Cohesion & Coherence
As we have seen, there appears to be an intimate link between parametricity, cohesion, and coherence, demonstrated (e.g.) by the above-given proof of induction for $S^{1}$. The existence of such a link is further supported by other recent developments in HoTT and related fields. E.g. Cavallo & Harper [5] have used their system of internal parametricity for Cubical Type Theory to derive coherence theorems for the smash product. More recently still, Kolomatskaia & Shulman [8] have introduced their system of Displayed Type Theory, which utilizes yet another form of internal parametricity to solve the previously-open problem of representing semi-simplicial types in HoTT.
In outlining the above framework for parametricity in terms of cohesion, I hope to have taken a first step toward the unification of these various systems that use parametricity to tackle instances of the coherence problem. Cavallo & Harper’s system is readily assimilated to this framework, as are other related systems that take a cubical approach to internal parametricity, such as that of Nuyts, Devriese & Vezzossi [11], since these all take their semantics in various topoi of bicubical sets, which all are sufficiently cohesive over corresponding topoi of cubical sets. On the other hand, Kolomatskaia & Shulman’s system cannot be assimilated in quite the same way, since their system takes its semantics in the $\infty$-topos of augmented semisimplicial spaces, which is not cohesive over spaces. It thus remains to be seen if the above framework can be generalized so as to be inclusive of this example as well. Alternatively, one might seek to generalize or modify Kolomatskaia & Shulman’s system so as to be interpretable in an $\infty$-topos that is cohesive over spaces, e.g. the $\infty$-topos of simplicial spaces. If this latter proves feasible, then this in turn would reveal yet further connections between parametricity via cohesion and another prominent attempt at a solution to the coherence problem, namely Riehl & Shulman’s Simplicial Type Theory, which takes its semantics in simplicial spaces [13].
@@
Acknowledgement
The origins of this paper trace to research I did as an undergraduate at Merton College, Oxford, in the Summer of 2020, as part of the Merton College Summer Projects Scheme. Naturally, this work was conducted at a time of considerable stress for myself and the world at large, and I am massively indebted to the academic support staff at Merton, who were immensely helpful in keeping me afloat at that time. I am particularly grateful to Katy Fifield, Jemma Underdown, and Jane Gover for the support they provided to me in the course of my undergraduate studies. More recently, I am grateful to Frank Pfenning and Steve Awodey for their encouragement in continuing to pursue this line of research.